Determine the mass defect and the binding energy in MeV/nucleon for 10146 Pd (atomic mass = 100.908287 amu). Proton = 1.00783 amu, neutron = 1.00866 amu, and 1 amu = 931.5 MeV?
Answer #1

The given nuclei is ^{101} _{46}Pd .

Hence, number of protons, n_p = 46

Number of neutrons, n_n = 101-46 = 55 .

Now, mass defect is defined as mass of the nuclei subtracted from the sum of mass of the protons and neutrons.

Hence,

\Delta m = \left (n_p \times m_p + n_n \times m_n \right ) - m_{Pd}

Where

m_p = mass \ of \ proton = 1.00783 \ amu

m_n = mass \ of \ neutron = 1.00866\ amu

m_{Pd}= 100.908287\ amu

Hence,

\Delta m = \left (n_p \times m_p + n_n \times m_n \right ) - m_{Pd} \\ \Rightarrow \Delta m = (46 \times 1.00783 \ amu + 55 \times 1.00866 \ amu) - 100.908287 \ amu \\ \Rightarrow \Delta m = (46.36018 \ amu + 55.4763 \ amu) - 100.908287 \ amu \\ \Rightarrow \Delta m = 0.928193 \ amu

Hence, the binding energy of the nucleus can be calculated by converting the amu units to MeV unit.

E = 0.928193 \ amu \times \frac{931.5 \ MeV}{1 \ amu} \approx 864.6 \ MeV

Now, binding energy per nucleon can be calculated by dividing the energy above by the number of nucleons i.e. 46+55=101.

Hence,

E_{per \ nucleon} = \frac{ 864.6 \ MeV}{101 \ nucleon} \approx {\color{Red} 8.560 \ MeV/nucleon}

Hence, the binding energy per nucleon of ^{101} _{46}Pd is about 8.560 MeV/nucleon.